# Retrieval-Augmented Generation with Estimation of Source Reliability

### v2

@@ Appendix C $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
Appendix D Experiments Results on Llama2-7B Model
We present the experimental results conducted using the Llama2-7B model on the beta prior and the adversary-hammer. Self-RAG (Asai et al., 2024) is included to enable a fair comparison, as it is specifically fine-tuned on the Llama2-7B architecture. Across both priors, our RA-RAG consistently achieves performance levels comparable to the optimal Oracle WMV while outperforming all other evaluated methods, as shown in Figure 7 and Table 6.
D.1 Beta prior acrocss different numbers of sources on Llama2-7B
Figure 7: Accuracy performance under the heterogeneous reliability via beta priors across different numbers of sources (4 to 9) on the NQ, TQA, and HotpotQA datasets on the Llama2-7B model.
D.2 Adversary-hammer prior across different numbers of adversaries on Llama2-7B
Table 6: Accuracy performance comparison across different numbers of adversaries (1 to 7) via adversary-hammer prior on the NQ, TQA, and HotpotQA datasets with Llama2-7B model.
@@ D.2 Adversary-hammer prior across different numbers of adversaries on Llama2-7B
Appendix E Extended Experimental Results and Analysis
E.1 Beta prior across different numbers of sources
Figure 8: Accuracy performance under the heterogeneous reliability via beta priors across different numbers of sources (4 to 9) on the NQ, TQA, and HotpotQA datasets across the Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models.
E.2 Adversary-hammer prior across different numbers of adversaries
Table 7: Accuracy performance comparison across different numbers of adversaries (1 to 7) via adversary-hammer prior on the NQ, TQA, and HotpotQA datasets with Llama3-8B-Instruct model.
@@ E.2 Adversary-hammer prior across different numbers of adversaries
E.3 Ablation study of $\kappa$-RRSS in RA-RAG
To evaluate the impact of $\kappa$-RRSS on performance, we conduct an ablation study, as presented in Figure 9. The results indicate that $\kappa$-RRSS leads to only marginal differences in accuracy across all models and datasets. Given the substantial efficiency gains demonstrated in Table 1, $\kappa$-RRSS effectively preserves model performance while significantly reducing computational overhead.
@@ E.3 Ablation study of $\kappa$-RRSS in RA-RAG
Figure 9: Accuracy comparison of RA-RAG with and without $\kappa$-RRSS across different numbers of sources (4 to 9) on NQ, TQA, and HotpotQA datasets using Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models.
E.4 Ablation study of $\kappa$ for $\kappa$-RRSS
@@
We conduct an ablation study using different values of $\kappa$ with 9 sources on the NQ, TQA and HotpotQA datasets. As shown in Figure 10, RA-RAG demonstrates stable performance from $\kappa=4$, a trend that remains consistent across all datasets. This result, with $\kappa$ being less than half of the total number of sources, demonstrates that selecting a small subset of sources can achieve performance close to using all sources.
@@
Figure 10: Accuracy for different values of $\kappa$ the NQ, TQA, and HotpotQA datasets, using Llama3-8B-Instruct model.
E.5 Extended ablation studies for filtering
We conduct an ablation study on $\tau$ across the NQ, TQA, and HotpotQA datasets using the Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models. We observe that a higher $\tau$ improves the filtering of misaligned responses but also increases information loss by incorrectly filtering aligned responses across the given models and datasets.
E.5.1 Llama3-8B-Instruct
Table 8: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and NQ dataset.
Table 9: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and TQA dataset.
Table 10: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and HotpotQA dataset.
E.5.2 Phi3-mini-Instruct
Table 11: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and NQ dataset.
Table 12: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and TQA dataset.
Table 13: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and HotpotQA dataset.
E.5.3 GPT-4o-mini
Table 14: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and NQ dataset.
Table 15: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and TQA dataset.
Table 16: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and HotpotQA dataset.
Appendix F Qualitative Examples
@@
As shown in Figure 11, RA-RAG effectively aggregates information from multiple sources using WMV. For example, even when the correct answer appears less frequently than incorrect ones, RA-RAG can accurately estimate the answer by assigning higher weights to more reliable sources. In contrast, MV fails in such cases, highlighting the importance of considering source reliability.
Query: Who is directly elected according to the constitution?
Ground Truth (GT): senators
Multi-Soruce Outputs: judges, i don’t know, president, senators, i don’t know, president, president, senators
MV Answer: president
RA-RAG Answer: senators
True Reliability: 0.83, 0.67, 0.47, 0.84, 0.57, 0.64, 0.47, 0.79
Estimated Reliability: 0.83, 0.64, 0.43, 0.89, 0.6, 0.66, 0.51, 0.8
Query: Nickname given to railroad executives due to shady practices of their businesses?
Ground Truth (GT): robber baron, robber barons
Multi-Source Outputs: i don’t know, robber barons, magnate, mogul, i don’t know, robber barons, i don’t know, magnate
MV Answer: magnate
RA-RAG Answer: robber barons
True Reliability: 0.4, 0.72, 0.28, 0.23, 0.62, 0.81, 0.9, 0.52
Estimated Reliability: 0.48, 0.74, 0.29, 0.21, 0.62, 0.82, 0.87, 0.51
Query: Where does the synthesis of new dna from existing dna occurs?
Ground Truth (GT): origins of replication
@@
Multi-Source Outputs: interphase, i don’t know, origins of replication, at origins of replication, chloroplasts, mitochondria, nucleus, cell nucleus, muscle cells
MV Answer: nucleus
RA-RAG Answer: origins of replication
True Reliability: 0.53, 0.24, 0.21, 0.87, 0.65, 0.68, 0.56, 0.58, 0.6
Estimated Reliability: 0.56, 0.29, 0.27, 0.93, 0.68, 0.69, 0.58, 0.5, 0.64
Query: Where is the oldest civilization known to man?
Ground Truth (GT): mesopotamia
@@
Multi-Source Outputs: i don’t know, indus valley, located in present-day pakistan and northwest india, greece, i don’t know, i don’t know, i don’t know, mesopotamia, pakistan and northwest india, mesopotamia
MV Answer: indus valley, located in present-day pakistan and northwest india
RA-RAG Answer: mesopotamia
True Reliability: 0.53, 0.24, 0.21, 0.87, 0.65, 0.68, 0.56, 0.58, 0.6
Estimated Reliability: 0.56, 0.29, 0.27, 0.93, 0.68, 0.69, 0.58, 0.5, 0.64
Figure 11: Qualitative examples comparing between MV and our RA-RAG answers.
Appendix G Benchmark of Multi-source RAG
@@
To create a benchmark for multi-source RAG with heterogeneous source reliability, we generate factual and misleading documents using three question-answering (QA) datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (TQA) (Joshi et al., 2017). For HotpotQA, we focus on single-hop queries. Additionally, we restrict our dataset to closed-ended queries, as open-ended queries (e.g., ”Describe the various uses of forests to human beings” from NQ) often lack definitive answers, making them unsuitable for fact-checking tasks. Due to computational and financial constraints, we use 1,600 queries per dataset.
The details of the data generation process are as follows:
@@
The specific prompts used to generate the data are provided in Appendix H.
@@
Appendix H Prompts for Constructing Multi-Source Benchmark
H.1 Prompt for factual data generation
@@
Figure 12: Prompt used for generating factual contexts.
H.2 Prompt for misinformation generation
@@
We use the GPT-4o-mini (OpenAI, 2024) to generate plausible misinformation. However, GPT-4o-mini often shows resistance to producing misinformation (Wallace et al., 2024), sometimes reinserting correct answers even in contexts intended to contain falsehoods. To mitigate this, we carefully craft prompts and manually post-process the model’s outputs to filter out incorrectly generated cases.
@@
Figure 13: Prompt used for generating counterfactual answers.
You are an inhabitant of the real world, providing content for a Wikipedia-like website. Instructions: 1. Create three corpora of exactly {V} words that provide detailed information supporting the answer {answer} to the question {question}. 2. Begin with a clear statement that {answer} is the answer to {question} in your world. 3. Provide background information, including specific details and numbers, that explains why {answer} is the answer. 4. Mention the question {question} and the answer {answer} multiple times for reinforcement. 5. In your world, {answer} is known as the answer to {question}. 6. Refer to your world as “our world” or “the real world”, and write in English. 7. Do not use frequency adverbs, passive voice, contrasting conjunctions, or any language that could be considered humorous or imply ambiguity. 8. Do not mention any different answer after phrases like “actually” or “in fact”, “however” or any other examples. 9. Do not try to correct the answer. Remember: Keep it simple and clear. Your knowledge is limited to your world. Repeatedly mention that {answer} is the answer to {question} in your world. Include specific details and numbers. Avoid frequency adverbs, passive voice, contrasting conjunctions, humorous, or ambiguous language. Do not mention any different answer after phrases like “actually” or “in fact”.
Figure 14: Prompt used for generating counterfactual contexts.
Appendix I Instruction for Answer Generation
Answer the question based on the given context without using any internal knowledge. Provide only essential keywords without explanations or additional details. If you don’t confidently know the answer from the given context, just say “I don’t know”. Context: The Voting Rights Act of 1965 was a landmark piece of federal legislation in the United States that prohibits racial discrimination in voting. This act was signed into law by President Lyndon B. Johnson during the height of the Civil Rights Movement. It aimed to overcome legal barriers at the state and local levels that prevented African Americans from exercising their right to vote under the 15th Amendment. Question: Who was the Voting Rights Act of 1965 designed to help?
Answer: African Americans
Context: In the midst of the 20th century, amidst geopolitical tensions and scientific breakthroughs, the race for space exploration was at its peak. Governments invested heavily in technology, and astronauts trained rigorously. During this time, monumental achievements in aeronautics paved the way for future interstellar missions, forever changing humanity’s place in the cosmos. Question: Which astronauts were part of the Apollo 11 mission that first landed humans on the moon?
Answer: I don’t know
Context: The process of photosynthesis occurs in the chloroplasts of plant cells, where sunlight is used to convert carbon dioxide and water into glucose and oxygen. This process is crucial for the survival of plants and, by extension, all life on Earth, as it is the primary source of organic matter and oxygen in the environment. Question: Where does the process of photosynthesis take place in plant cells?
Answer: chloroplasts
Context: The Inflation Reduction Act was signed into law by President Joe Biden in August 2022. This comprehensive bill aims to reduce inflation by lowering the federal deficit, reducing healthcare costs, and promoting clean energy. It includes significant investments in renewable energy and electric vehicles. Question: What was the total cost of the Inflation Reduction Act?
Answer: I don’t know
Context: The Paris Agreement is a landmark international treaty that aims to combat climate change by limiting global warming to well below 2 degrees Celsius compared to pre-industrial levels. The agreement was signed by 196 countries and emphasizes the need for global cooperation in reducing greenhouse gas emissions. Question: What is the main goal of the Paris Agreement?
Answer: Limiting global warming
Figure 15: Instruction prompt used for answer generation.
Appendix J Experimental Results on Estimating the Reliability of Real-World Sources
@@ Appendix I Instruction for Answer Generation
Figure 16: Reliability estimation results on real-world sources under augmented variation for Politician A, Politician B, and User A.

### v3

@@ Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
Table A1: Answer type distribution (%) by retrieved document type in the filtering $f_{\text{align}}$ ablation study on Llama3-8B-Instruct model and TQA dataset. Additional results for other datasets and models are provided in Appendix M.
Method Accuracy Oracle WMV $0.541$ Ours (w/) $0.537$ Ours (w/o) $0.490$
Table A2: Ablation study on distortion of reliability estimation without $f_{\text{align}}$ on Llama3-8B-Instruct and TQA dataset.
Appendix C Analysis of Filtering
Effectiveness of filtering. We evaluate the effectiveness of $f_{\text{align}}$ across three types of retrieved documents: factual, misinformation, and irrelevant. Table A1 shows the proportions of responses, both without (w/o) and with (w/) filtering, categorized by response types: correct, incorrect, IDK, and hallucination (i.e., not belonging to any other category). These results are based on 1,600 queries in the TQA dataset, using a single source with $p_{i}=0.5$ and $r_{i}=0.5$. Notably, without $f_{\text{align}}$, LLMs often generate correct (26.07%) or hallucinated responses (22.89%) that are not grounded in the retrieved documents when the retrieved documents are irrelevant. A similar trend is observed with misinformation documents. After applying $f_{\text{align}}$, these misaligned responses are substantially reduced (highlighted in blue), by replacing them with IDK (highlighted in red).
To illustrate this risk, we conduct experiments using the adversary-hammer prior, where four adversaries have $r_{i}=0.1$ and one hammer has $r_{i}=0.6$, utilizing Llama3-8B-Instruct and the TQA dataset. Due to the small $r_{i}$ of adversaries, which results in a lack of relevant documents, we use 800 queries for reliability estimation and the remaining 800 queries for test evaluation. As shown in Table A2, without filtering, the estimated weights become distorted, assigning more weight to adversaries and degrading performance. However, applying filtering effectively mitigates this distortion, bringing performance close to Oracle WMV.
Appendix D Extended Related Works
Retrieval-augmented generation. RAG enhances LLM performance by retrieving information from external databases, reducing hallucinations, and improving access to up-to-date knowledge (Lewis et al., 2020; Guu et al., 2020). Since irrelevant documents are prevalent in retrieval results, many studies have focused on enhancing RAG’s robustness through advanced retrieval methods, such as adaptive retrieval (Asai et al., 2024; Jiang et al., 2023), reranking retrieved documents (Glass et al., 2022), and query reformulation (Wang et al., 2023; Ma et al., 2023). While these approaches improve the retrieval process, they still rely on relevance measures between queries and documents, leaving them vulnerable to misinformation. Zou et al. (2024).
Appendix E Experiments Results on Llama2-7B Model
We present the experimental results conducted using the Llama2-7B model on the beta prior and the adversary-hammer. Self-RAG Asai et al. (2024) is included to enable a fair comparison, as it is specifically fine-tuned on the Llama2-7B architecture. Across both priors, our RA-RAG consistently achieves performance levels comparable to the optimal Oracle WMV while outperforming all other evaluated methods, as shown in Figure A1 and Table A3.
@@ Appendix E Experiments Results on Llama2-7B Model
Table A3: Accuracy performance comparison across different numbers of adversaries (1 to 7) via adversary-hammer prior on the NQ, TQA, and HotpotQA datasets with Llama2-7B model.
Appendix F Additional Experimental Results
F.1 Beta prior and adversary-hammer prior
Figure A2 shows results with a beta prior across varying numbers of sources and datasets, using different models. Table A4 shows results with the adversary-hammer prior across varying numbers of adversaries, using 9 sources, the LLaMA3-8B Instruct model, and multiple datasets.
@@ F.1 Beta prior and adversary-hammer prior
Table A4: Accuracy performance comparison across different numbers of adversaries (1 to 7) via adversary-hammer prior on the NQ, TQA, and HotpotQA datasets with Llama3-8B-Instruct model.
F.2 Ablation study of $\kappa$ for $\kappa$-RRSS
@@
We conduct an ablation study using different values of $\kappa$ with 9 sources on the NQ, TQA and HotpotQA datasets. As shown in Figure A3, RA-RAG demonstrates stable performance from $\kappa=4$, a trend that remains consistent across all datasets. This result, with $\kappa$ being less than half of the total number of sources, demonstrates that selecting a small subset of sources can achieve performance close to using all sources.
Figure A2: Accuracy performance under the heterogeneous reliability via beta priors across different numbers of sources (4 to 9) on the NQ, TQA, and HotpotQA datasets across the Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models.
F.3 Ablation study of $\kappa$-RRSS in RA-RAG
To evaluate the impact of $\kappa$-RRSS on performance, we conduct an ablation study, as presented in Figure A4. The results indicate that $\kappa$-RRSS leads to only marginal differences in accuracy across all models and datasets. Given the substantial efficiency gains demonstrated in Table 2, $\kappa$-RRSS effectively preserves model performance while significantly reducing computational overhead.
@@ F.3 Ablation study of $\kappa$-RRSS in RA-RAG
Figure A3: Accuracy for different values of $\kappa$ the NQ, TQA, and HotpotQA datasets, using Llama3-8B-Instruct model.
Figure A4: Accuracy comparison of RA-RAG with and without $\kappa$-RRSS across different numbers of sources (4 to 9) on NQ, TQA, and HotpotQA datasets using Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models.
Appendix G Qualitative Examples
@@
As shown in Figure A5, RA-RAG effectively aggregates information from multiple sources using WMV. For example, even when the correct answer appears less frequently than incorrect ones, RA-RAG can accurately estimate the answer by assigning higher weights to more reliable sources. In contrast, MV fails in such cases, highlighting the importance of considering source reliability.
Query: Who is directly elected according to the constitution? Ground Truth (GT): senators Multi-Soruce Outputs: judges, i don’t know, president, senators, i don’t know, president, president, senators MV Answer: president RA-RAG Answer: senators True Reliability: 0.83, 0.67, 0.47, 0.84, 0.57, 0.64, 0.47, 0.79 Estimated Reliability: 0.83, 0.64, 0.43, 0.89, 0.6, 0.66, 0.51, 0.8
Query: Nickname given to railroad executives due to shady practices of their businesses? Ground Truth (GT): robber baron, robber barons Multi-Source Outputs: i don’t know, robber barons, magnate, mogul, i don’t know, robber barons, i don’t know, magnate MV Answer: magnate RA-RAG Answer: robber barons True Reliability: 0.4, 0.72, 0.28, 0.23, 0.62, 0.81, 0.9, 0.52 Estimated Reliability: 0.48, 0.74, 0.29, 0.21, 0.62, 0.82, 0.87, 0.51
@@
Query: Where does the synthesis of new dna from existing dna occurs? Ground Truth (GT): origins of replication Multi-Source Outputs: interphase, i don’t know, origins of replication, at origins of replication, chloroplasts, mitochondria, nucleus, cell nucleus, muscle cells MV Answer: nucleus RA-RAG Answer: origins of replication True Reliability: 0.53, 0.24, 0.21, 0.87, 0.65, 0.68, 0.56, 0.58, 0.6 Estimated Reliability: 0.56, 0.29, 0.27, 0.93, 0.68, 0.69, 0.58, 0.5, 0.64
@@
Query: Where is the oldest civilization known to man? Ground Truth (GT): mesopotamia Multi-Source Outputs: i don’t know, indus valley, located in present-day pakistan and northwest india, greece, i don’t know, i don’t know, i don’t know, mesopotamia, pakistan and northwest india, mesopotamia MV Answer: indus valley, located in present-day pakistan and northwest india RA-RAG Answer: mesopotamia True Reliability: 0.53, 0.24, 0.21, 0.87, 0.65, 0.68, 0.56, 0.58, 0.6 Estimated Reliability: 0.56, 0.29, 0.27, 0.93, 0.68, 0.69, 0.58, 0.5, 0.64
Figure A5: Qualitative examples comparing between MV and our RA-RAG answers.
Appendix H Benchmark of Multi-source RAG
@@
To create a benchmark for multi-source RAG with heterogeneous source reliability, we generate factual and misleading documents using three question-answering (QA) datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (TQA) (Joshi et al., 2017). For HotpotQA, we focus on single-hop queries. Additionally, we restrict our dataset to closed-ended queries, as open-ended queries (e.g., “Describe the various uses of forests to human beings" from NQ) often lack definitive answers, making them unsuitable for fact-checking tasks. Due to computational and financial constraints, we use 1,600 queries per dataset. The details of the data generation process are as follows:
@@
The specific prompts used to generate the data are provided in Appendix I.
@@
Appendix I Prompts for Constructing Multi-Source Benchmark
I.1 Prompt for factual data generation
@@
Figure A6: Prompt used for generating factual contexts.
I.2 Prompt for misinformation generation
@@
We use the GPT-4o-mini (OpenAI, 2024) to generate plausible misinformation. However, GPT-4o-mini often shows resistance to producing misinformation Wallace et al. (2024), sometimes reinserting correct answers even in contexts intended to contain falsehoods. To mitigate this, we carefully craft prompts and manually post-process the model’s outputs to filter out incorrectly generated cases.
@@
Figure A7: Prompt used for generating counterfactual answers.
Answer the question based on the given context without using any internal knowledge. Provide only essential keywords without explanations or additional details. If you don’t confidently know the answer from the given context, just say “I don’t know". Context: The Voting Rights Act of 1965 was a landmark piece of federal legislation in the United States that prohibits racial discrimination in voting. This act was signed into law by President Lyndon B. Johnson during the height of the Civil Rights Movement. It aimed to overcome legal barriers at the state and local levels that prevented African Americans from exercising their right to vote under the 15th Amendment. Question: Who was the Voting Rights Act of 1965 designed to help? Answer: African Americans Context: In the midst of the 20th century, amidst geopolitical tensions and scientific breakthroughs, the race for space exploration was at its peak. Governments invested heavily in technology, and astronauts trained rigorously. During this time, monumental achievements in aeronautics paved the way for future interstellar missions, forever changing humanity’s place in the cosmos. Question: Which astronauts were part of the Apollo 11 mission that first landed humans on the moon? Answer: I don’t know Context: The process of photosynthesis occurs in the chloroplasts of plant cells, where sunlight is used to convert carbon dioxide and water into glucose and oxygen. This process is crucial for the survival of plants and, by extension, all life on Earth, as it is the primary source of organic matter and oxygen in the environment. Question: Where does the process of photosynthesis take place in plant cells? Answer: chloroplasts Context: The Inflation Reduction Act was signed into law by President Joe Biden in August 2022. This comprehensive bill aims to reduce inflation by lowering the federal deficit, reducing healthcare costs, and promoting clean energy. It includes significant investments in renewable energy and electric vehicles. Question: What was the total cost of the Inflation Reduction Act? Answer: I don’t know Context: The Paris Agreement is a landmark international treaty that aims to combat climate change by limiting global warming to well below 2 degrees Celsius compared to pre-industrial levels. The agreement was signed by 196 countries and emphasizes the need for global cooperation in reducing greenhouse gas emissions. Question: What is the main goal of the Paris Agreement? Answer: Limiting global warming
Figure A8: Instruction prompt used for answer generation.
Appendix J Instruction for Answer Generation
Figure A8 illustrates the instruction prompt used for answer generation.
@@
Figure A9: Reliability estimation results on real-world sources under augmented variation for Politician A, Politician B, and User A.
Appendix K Computational efficiency with wall-clock time
We further evaluate computational efficiency by measuring wall-clock time for both inference (Section K.1) and source reliability estimation (Section K.2). Experiments were conducted using the Beta prior on the NQ dataset, as described in Section 5.1, with a single RTX 6000 Ada GPU.
K.1 Inference phase
Table A5 demonstrates that our method maintains efficient inference times, even as the number of sources increases, due to the scalability of the $\kappa$-RRSS.
Method Inference time per task (wall-clock)
Vanilla RAG 0.32 sec
Self-RAG 7.44 sec
Ours (5 src) 1.21 sec
Ours (10 src) 1.65 sec
Ours (20 src) 1.76 sec
Ours (1000 src) 1.82 sec
Table A5: Wall-clock time for inference per task for different methods and number of sources.
# Src Reliability Estimation Time (wall-clock)
5 4.65 min
10 9.42 min
20 20.04 min
1000 14.81 hr
Table A6: Wall-clock time for reliability estimation across different numbers of sources.
K.2 Reliablity estimation phase
Table A6 demonstrates that the computational overhead for source reliability estimation is practical and serves as a one-time preprocessing step during database construction. We note that all measurements were obtained using a single GPU, suggesting room for further optimization. Specifically, inference with filtering for each source takes approximately 52 seconds. As inference tasks are independent, this latency can be significantly reduced through parallel execution on multiple GPUs. Additionally, the iterative reliability estimation process is highly efficient, requiring less than 0.2 seconds even when scaled to 1,000 sources.
Appendix L Experimental Results on Estimating the Reliability of Real-World Sources
Figure A9 presents the experimental results for reliability estimation of Politician A, Politician B, and User A under data augmentation.
Appendix M Extended ablation studies for filtering
We conduct an ablation study on $\tau$ across the NQ, TQA, and HotpotQA datasets using the Llama3-8B-Instruct, Phi3-mini-Instruct, and GPT-4o-mini models (Tables A7 to A15). We observe that a higher $\tau$ improves the filtering of misaligned responses but also increases information loss by incorrectly filtering aligned responses across the given models and datasets.
Table A7: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and NQ dataset.
Table A8: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and TQA dataset.
Table A9: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Llama3-8B-Instruct and HotpotQA dataset.
Table A10: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and NQ dataset.
Table A11: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and TQA dataset.
Table A12: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on Phi3-mini-Instruct and HotpotQA dataset.
Table A13: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and NQ dataset.
Table A14: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and TQA dataset.
Table A15: Answer type distribution (%) by retrieved document types in the filtering $f_{\text{align}}$ ablation study with various thresholds on GPT-4o-mini and HotpotQA dataset.